# Adaptive Evidence Budgeting for Scalable Long-Document Reranking with LLMs

### v4

@@ 6.1 Factorial Ablation: AEB and SA
We conduct a $\times2$ factorial ablation on DL19 to isolate the effects of (i) Adaptive Evidence Budgeting (AEB) (dynamic stopping) and (ii) Summary Augmentation (SA). All runs use the same reranker, candidate set, budget cap ($p_{\max}{=}600$), score normalization (minmax, T=1.0T{=}1.0), and random seed; we only toggle these two components.
@@
Table 5: $\times2$ factorial ablation on DL19 (cross selector). Same reranker, candidates, budget ($p_{\max}{=}600$), normalization (minmax, T=1.0T{=}1.0), and seed; only AEB and SA are toggled.
@@ 6.2 Ablation on Summary Augmentation
Variant Budget nDCG@10
EviRerank (SA, default) 480 + sum 0.739
Random-3-block summary 480 + sum 0.731
EviRerank (no SA), longer input 600 (no sum) 0.728
Table 6: Summary augmentation ablation on DL19 under the same total cap ($p_{\max}{=}600$) (fixed setting).
We test whether SA can be replaced under the same cap ($p_{\max}=600$) by fixing the non-AEB setting and comparing SA (480+sum) against Random-3-Block (480+sum) and a longer-input no-summary control (600) (Table 6). SA performs best, and even random-block cues outperform spending the same budget on longer query-focused blocks, suggesting that compact summary-style cues add complementary signals beyond simply increasing input length.
@@ 7 Conclusion
We presented EviRerank, an evidence-guided framework for long-document reranking that selects salient blocks under a strict token budget and scores the resulting evidence context with a decoder-only LLM. EviRerank introduces Adaptive Evidence Budgeting (AEB) to pack evidence query-adaptively via early stopping, improving the accuracy–efficiency trade-off under the same budget cap. We also studied an optional Summary Augmentation (SA) module; controlled ablations show that its gains are not merely from longer inputs. Across TREC DL’19, DL’23, and MLDR-zh, EviRerank consistently outperforms full-document LLM reranking, pooling variants, and strong block-selection baselines. Overall, the results suggest a practical takeaway: for long-document reranking, how to allocate the evidence budget is as important as which evidence to include.
@@
First, our SA module is lightweight and query-agnostic; although it is budget-friendly and works well in practice, richer cues (e.g., query-conditioned or multi-facet summaries) may be beneficial for documents with diverse intents. Second, our evaluation focuses on three long-document reranking benchmarks and one decoder-only reranker family; broader coverage (more domains and different reranker architectures) would strengthen generalization claims, which would be future work.
@@ Appendix A Training Memory Comparison
Figure 4 reports peak GPU memory usage during training on an NVIDIA H20 (96 GB). Full-document RankLLaMA, which consumes up to 4096 tokens per query–document pair, exceeds the 96 GB capacity even with batch size 1; we therefore mark it as >96>96 GB in the figure. This highlights the high memory footprint of full-context training under long inputs.
In contrast, EviRerank trains within a much smaller memory budget by constructing a compact evidence context under a strict cap. Across selectors, peak memory stays around 22–23 GB.
Pooling-based baselines (RankLLaMA-MaxP/AvgP) are the most memory-efficient (17,235/17,233 MB), but they are substantially less effective than EviRerank (Sec. 5.1). Overall, budget-aware evidence construction offers a practical accuracy–resource trade-off for long-document reranking.
@@ B.1 Attention Heatmaps of Specific Examples
To shed light on this and answer RQ1, we propose first analyzing the attention heatmaps of the RankLLaMA model with several specific examples, which has been fine-tuned on the MSMARCO passage ranking dataset222available at: https://huggingface.co/castorini/rankLlama-v1-7b-lora-passage. Clark et al. (2019) propose a similar investigation for BERT and find a significant portion of BERT’s attention is focused on the delimiter token, and certain attention heads align well with linguistic features such as syntax and coreference. However, recent LLMs are uni-directional and decoder-only, especially for the IR focused LLMs, and may display a different behavior regarding how tokens attend to each other. We used the BertViz tool Vig (2019) to explore these attention patterns in the decoder-only LLM RankLLaMA when processing various query-passage pairs.
@@
    In contrast, head 23 in layer 1 (Fig. 6(e)) continues to exhibit broad attention across all tokens. Although the attention may be weak, it reveals that the head does not clearly distinguish between relevant and irrelevant content. Similarly, head 24 in layer 31 (Fig. LABEL:fig.4e) shows that the token “Paris” in the document attends not only to the query token “Paris” but also weakly to irrelevant tokens such as “apple,” “is,” and “a.”
@@ B.3.1 Metric Definitions
    $\mathbf{a}$ $=\big(\mathrm{Attn}(C_{i})\big)_{i=1}^{M},\quad\mathbf{r}=\big(\mathrm{Rel}(C_{i})\big)_{i=1}^{M}.$
@@ B.3.2 Experimental Setup
We conduct controlled experiments on 500 query-relevant document pairs sampled from the same development set (MS MARCO) as Section B.2. Each document is truncated to 1200 tokens, and segmented into 64-token chunks for analysis. The ARAS and PCR metrics are then computed across three representative attention heads previously identified: Layer 1 Head 23, Layer 8 Head 25, and Layer 31 Head 24.
@@ B.4 Summary: Implications for Long Document Retrieval
(a) Insert noise document texts before relevant document
(b) Insert noise document texts after relevant document
(c) Insert noise document texts before relevant document
(d) Insert noise document texts after relevant document
(e) Insert noise document texts before relevant document
(f) Insert noise document texts after relevant document

### v5

@@ 6.1 Factorial Ablation: AEB and SA
We conduct a $\times2$ factorial ablation on DL19 to isolate the effects of (i) AEB (dynamic stopping) and (ii) SA. We enforce $p_{\max}{=}600$ as a hard upper bound on doc-side content; the realized length can be below the cap because blocks are packed atomically (we do not force filling the remaining budget by splitting a block). All runs use the same reranker, candidate set, minmax normalization, and random seed; we only toggle these two components.
@@
Table 5: $\times2$ factorial ablation on DL19 (cross selector).
@@ 6.2 Ablation on Summary Augmentation
Variant nDCG@10
EviRerank (SA, default) 0.739
Random blocks summary 0.731
EviRerank (no SA) 0.728
Table 6: Summary augmentation ablation on DL19 under the same max cap.
We evaluate whether SA can be replaced under the same cap ($p_{\max}{=}600$) by fixing the non-AEB setting and comparing SA against (i) a random same block count summary and (ii) a no-summary control (using the full cap for evidence blocks) (Table 6). SA performs best, showing that adding our SA cue under the same cap improves effectiveness beyond allocating the entire budget to additional query-focused blocks or using a random-block cue.
@@ 7 Conclusion
We presented EviRerank, an evidence-guided framework for long-document reranking that selects salient blocks under a hard token cap and scores the resulting evidence context with a decoder-only LLM. EviRerank introduces Adaptive Evidence Budgeting (AEB) to pack evidence query-adaptively via early stopping, improving the accuracy–efficiency trade-off under the same cap. We also studied Summary Augmentation (SA) and show via controlled ablations that its gains are not merely from longer inputs. Across TREC DL’19, DL’23, and MLDR-zh, EviRerank consistently outperforms full-document LLM reranking, pooling variants, and strong block-selection baselines. Overall, EviRerank shows that even though decoder-only LLM rerankers are powerful, effective long-document reranking hinges on evidence construction under a fixed input cap: deciding not only which blocks to select, but also how much evidence to include and whether to add a global summary cue.
@@
First, our SA module is lightweight and query-agnostic; although it is budget-friendly and works well in practice, richer cues (e.g., multi-facet summaries) may be beneficial for documents with diverse intents. Second, our evaluation focuses on three long-document reranking benchmarks and one decoder-only reranker family; broader coverage (more domains and different reranker architectures) would strengthen generalization claims, which would be future work.
@@ Appendix A Training Memory Comparison
Fig. 4 reports peak GPU memory usage during training on an NVIDIA H20 (96 GB). Full-document RankLLaMA, which consumes up to 4096 tokens per query-document pair, exceeds the 96 GB capacity even with batch size 1; we therefore mark it as >96>96 GB in the figure. This highlights the high memory footprint of full-context training under long inputs.
In contrast, EviRerank trains within a much smaller memory budget by constructing a compact evidence context under a strict cap. Across selectors, peak memory stays around 22-23 GB.
Pooling-based baselines (RankLLaMA-MaxP/AvgP) are the most memory-efficient (17,235/17,233 MB), but they are substantially less effective than EviRerank (Sec. 5.1). Overall, budget-aware evidence construction offers a practical accuracy-resource trade-off for long-document reranking.
@@ B.1 Attention Heatmaps of Specific Examples
To shed light on this and answer RQ1, we propose first analyzing the attention heatmaps of the RankLLaMA model with several specific examples, which has been fine-tuned on the MSMARCO passage ranking dataset333available at: https://huggingface.co/castorini/rankLlama-v1-7b-lora-passage. Clark et al. (2019) propose a similar investigation for BERT and find a significant portion of BERT’s attention is focused on the delimiter token, and certain attention heads align well with linguistic features such as syntax and coreference. However, recent LLMs are uni-directional and decoder-only, especially for the IR focused LLMs, and may display a different behavior regarding how tokens attend to each other. We used the BertViz tool Vig (2019) to explore these attention patterns in the decoder-only LLM RankLLaMA when processing various query-passage pairs.
@@
    In contrast, head 23 in layer 1 (Fig. 6(e)) continues to exhibit broad attention across all tokens. Although the attention may be weak, it reveals that the head does not clearly distinguish between relevant and irrelevant content. Similarly, head 24 in layer 31 (Fig. 8(e)) shows that the token “Paris” in the document attends not only to the query token “Paris” but also weakly to irrelevant tokens such as “apple,” “is,” and “a.”
@@ B.3.1 Metric Definitions
    $\mathbf{a}$ $=\big(\mathrm{Attn}(C_{i})\big)_{i=1}^{M},$
    $\mathbf{r}$ $=\big(\mathrm{Rel}(C_{i})\big)_{i=1}^{M}.$
@@ B.3.2 Experimental Setup
We conduct controlled experiments on 500 query-relevant document pairs sampled from the same development set (MS MARCO) as Section B.2. Each document is truncated to 1200 tokens, and segmented into 64-token chunks for analysis. The ARAS and PCR metrics are then computed across three representative attention heads previously identified: Layer 1 Head 23, Layer 8 Head 25, and Layer 31 Head 24. The figures are shown in Fig. 10.
@@ B.4 Summary: Implications for Long Document Retrieval
(a) ARAS of Layer 1 Head 23
(b) PCR of Layer 1 Head 23
(c) ARAS of Layer 8 Head 25
(d) PCR of Layer 8 Head 25
(e) ARAS of Layer 31 Head 24
(f) PCR of Layer 31 Head 24